At a certain gas station, 40% of the customers use regular unleaded gas, 35 % use extra unleaded gas, and the remainder use premium unleaded gas. Of those using regular gas, only 42% fill their tanks. Of those customers using extra unleaded, 34% fill their tanks, whereas of those using premium unleaded, 28% fill their tanks. What is the probability that the next customer will fill their tank?
Answer #1

P(Customer use premium gas) = 1 - (0.40 + 0.35) = 0.25

Hence,

P(Next customer will fill their tank)

= 0.40*0.42 + 0.35*0.34 + 0.25*0.28

= 0.3570